O KINETICS AND EQUILIBRIUM Deducing a rate law from initial reaction rate data Some measurements of the initial rate of a cer please help me solve this, i only get ONE try so please please make sure the answer is right, thanks!!
Answer #1

The answer in three significant figure is 17.4.

Please rate the answer, if you understand it. Thanks.

H2+ J2 2HI Ralt= k[a 40 data & data ( (os pi R,= 53-0 MIS ; H].-1-28M;エ]= 2-30M C R2= 6-54 M)S ; Hl- 1-28 M; [エ]= 0-284-M 4Tdata & date 3) . Ee T2-30M 2.30M H21-28 M RI= 53.0 MIS R3= 5.38 MIS D 130 M H2 kf12312-3079 Rfo-130T 2-3078 RI R3 53.0 5.38 9